How many ways can 6 people sit together in a row
a) If two people sit together?
b) How many ways can 6 people sit in a circle of two people sit together?

Please explain my second homework question to me. Thank you!

Found 2 solutions by math_helper, ikleyn:
Answer by math_helper(2461)


With no restrictions, 6 people can sit in 6! = 720 ways
Having two people sit together means we can treat those two as a unit, and
we get effectively 5 people to be arranged: 5! = 120 ways
BUT we must multiply this by 2! = 2 because the two-person unit can be formed in 2! ways (AB and BA): 2*120 = 240.
a) 240 ways


b) For this part, the physical configuration changes things. It is like the row A-B-C-D-E-F but curved around so F is actually understood to be next to A.
I'll try to draw it:

B - C
/ \
A D
\ /
F - E
This greatly reduces the number of permutations (seating arrangements). Here we assume it doesn't matter if, say, F is at "7 O'clock" --- if we rotated the above arrangement so F was at, say, "11 O'clock" but kept all the relative positions intact, it would still be the same arrangement.

The process:
Seat A anywhere. 'A' serves as a reference point in a way, leaving 5 seats for the others to occupy around A. Those 5 others can be arranged in 5! = 120 ways. And that is all of the ways they can be seated.
b) 120 ways
------
EDIT: Oh sorry, I think you wanted the circle configuration for two people sitting together (say AB, BA)... In this case it is like before: you effectively have 5 people (with one "person" a two-person "unit") so you have 4! = 24 ways to arrange those, times 2! = 2 ways to arrange the two-person unit (AB or BA). That's 2*24 = 48 ways.


Answer by ikleyn(52778)

.
How many ways can 6 people sit together in a row
a) If two people sit together?
b) How many ways can 6 people sit in a circle of two people sit together?

Please explain my second homework question to me. Thank you!
~~~~~~~~~~~~~

In part b), the correct  ANSWER  is  2*4! = 2*24 = 48 different possible ways.


Again, we consider these two persons as one unit and place them as a reference unit.


Then we order/permute the remaining  6 - 2 = 4 persons in 4! = 24 ways.


We multiply the number 24 by two, since there are two way to order that two persons, who are sitting together.

------------

To see many other similar  (and different)  solved problems,  look into the lesson
    - Persons sitting around a cicular table
in this site,  and learn the subject from there.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Combinatorics: Combinations and permutations".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.